# Structured Kolmogorov-Arnold Neural ODEs for Interpretable Learning and Symbolic Discovery of Nonlinear Dynamics

### v1

@@ 4.2 Van der Pol Oscillator
Figure 3: Results on the Van der Pol oscillator. Top left: Symbolic governing equation discovered by $\text{KAN}_{\text{symbolic}}$; a node receiving displacement input $x$ clearly exhibits the expected quadratic nonlinear relationship. Top right: Predicted system observables (accelerations) using the SKANODE framework. SKANODE achieves high accuracy in reconstructing accelerations, directly from latent dynamics without requiring a separate observation model. Bottom: Inferred latent state variables compared to ground truth. The latent states recovered by SKANODE correspond directly to physically meaningful displacement and velocity trajectories, in contrast to ANODE and SONODE, whose latent states lack clear physical interpretation and serve primarily as abstract representations for reconstructing observables.
We compare the performance of SKANODE, ANODE, and SONODE on this system. The symbolic governing equation extracted by $\text{KAN}_{\text{symbolic}}$ within SKANODE is shown in Figure 3, where the identified expression accurately reflects the true system dynamics. The symbolic network clearly reveals a distinct quadratic node that receives the displacement input $x$ and multiplies with the velocity input $\dot{x}$, consistent with the nonlinear damping characteristic of the Van der Pol oscillator. The corresponding predicted accelerations and recovered latent states are also presented. SKANODE accurately reconstructs both the observable accelerations and the latent states, recovering physically meaningful displacement and velocity trajectories. In contrast, ANODE fails to capture the system dynamics entirely, producing large deviations even in reconstructing the measured accelerations. SONODE performs better than ANODE in predicting the observables but still exhibits notable discrepancies and does not recover physically interpretable latent states. Its latent representations remain abstract, as the model lacks explicit physical structure. These results highlight the advantage of SKANODE in simultaneously achieving high predictive accuracy and interpretable latent dynamics through its structured state-space formulation and symbolic discovery capability.
@@ 4.3 F-16 Aircraft
The last example expands to a real-world complex system. The study presented by [62] details the experimental data collected from a full-scale F-16 aircraft during a ground vibration test master class. To simulate real-world conditions, two dummy payloads were mounted on the wing tips of the aircraft. The setup includes 145 acceleration sensors distributed across the aircraft, under excitation of a shaker installed beneath the right wing to apply multisine excitations over a frequency range of 2 to 15 Hz. Data from three specific sensors were made public: one at the excitation point, one on the right wing near the nonlinear interface of interest, and one on the payload adjacent to the same interface. These measurements were taken at a sampling frequency of 400 Hz. The interfaces, consisting of T-shaped connectors on the payload side, slid through a rail attached to the wing side, were identified by preliminary investigation as primary sources of nonlinearity in the aircraft’s structural dynamics, particularly at the right-wing-to-payload interface.
@@
Building on these observations, we extend SKANODE by introducing an additional latent hysteretic state $a_{h}$, leading to the following structured formulation:
@@
where $h_{\theta}$ captures the hysteretic dynamics and $\alpha\in[0,1]$ weights the relative contributions of elastic stiffness and hysteretic components. Through this extended structured state-space design, SKANODE is able to identify a symbolic governing equation, as presented in Figure 5. The resulting equation again aligns with the findings of [63], confirming that under low-level excitation, the system behavior is dominated by linear stiffness with mild hysteretic effects. This example highlights the flexibility of SKANODE in capturing real-world nonlinear dynamics through structured latent state design and symbolic equation discovery.
@@ 5 Conclusion
In this work, we propose Structured Kolmogorov–Arnold Neural ODEs (SKANODE), a novel framework that integrates structured state-space modeling with symbolic equation discovery for learning interpretable dynamics from partially observed nonlinear systems. By embedding a physically meaningful latent state structure and leveraging the Kolmogorov–Arnold Network (KAN) as both a universal function approximator and symbolic learner, SKANODE bridges the gap between the expressive capacity of deep learning and the interpretability required for scientific modeling. This approach enables the direct recovery of physically meaningful latent states, such as displacements and velocities, while simultaneously extracting explicit symbolic governing equations directly from indirect measurements like accelerations.
Comprehensive experiments on multiple nonlinear dynamical systems demonstrate that SKANODE outperforms existing NODE variants in both predictive accuracy and interpretability. These results highlight the importance of combining inductive structural biases with symbolic discovery capabilities to guide the learning process toward solutions that are physically consistent, robust, and interpretable. While the structured state-space formulation is highly effective for a broad class of regular dynamical systems, it may limit adaptability when applied to systems exhibiting more complex hybrid behaviors, discontinuities, or non-smooth switching dynamics that deviate from classical formulations. In such cases, overly rigid structure may risk overfitting to spurious regularities or failing to capture important unmodeled phenomena. Addressing such challenges may require extending SKANODE with more flexible hybrid formulations, adaptive state-space decompositions, or learnable switching mechanisms that can preserve physical interpretability while allowing for richer dynamic behaviors. The ability to uncover governing equations from indirect observations makes SKANODE particularly well-suited for scientific and engineering domains where full-state measurements are often impractical, and model transparency is critical for trustworthy decision-making. Future work will explore the integration of domain-specific physics priors and constraints within the symbolic extraction process.
@@ Appendix A SKANODE Recovers True System Dynamics
In this section, we theoretically demonstrate that, under certain conditions and assuming idealized loss minimization in the theoretical setting, the proposed Structured Kolmogorov–Arnold Neural ODE (SKANODE) framework is capable of exactly recovering the true governing dynamics.

### v2

@@ 4.2 Van der Pol Oscillator
Figure 3: Results on the Van der Pol oscillator. Top left: Symbolic governing equation discovered by $\text{KAN}_{\text{symbolic}}$, compared against the numerical baseline. The identified node receiving displacement input $x$ correctly captures the expected quadratic nonlinearity. Top right: Predicted system observables (accelerations) obtained using the SKANODE framework. SKANODE accurately reconstructs accelerations directly from the inferred latent dynamics, without the need for a separate observation model. Bottom: Inferred latent state variables compared to ground truth. The latent states recovered by SKANODE correspond closely to physically meaningful displacement and velocity trajectories, unlike ANODE and SONODE, whose latent states remain abstract representations without clear physical interpretation.
The performance of SKANODE, ANODE, and SONODE is compared on this system. The symbolic governing equation extracted by $\text{KAN}_{\text{symbolic}}$ within SKANODE is shown in Figure 3, where the identified expression accurately reflects the true system dynamics, outperforming the numerical baseline. The symbolic network clearly reveals a distinct quadratic node that receives the displacement input $x$ and multiplies with the velocity input $\dot{x}$, consistent with the nonlinear damping characteristic of the Van der Pol oscillator. The corresponding predicted accelerations and recovered latent states are also presented. SKANODE accurately reconstructs both the observable accelerations and the latent states, recovering physically meaningful displacement and velocity trajectories. In contrast, ANODE fails to capture the system dynamics entirely, producing large deviations even in reconstructing the measured accelerations. SONODE performs better than ANODE in predicting the observables but still exhibits notable discrepancies and does not recover physically interpretable latent states. Its latent representations remain abstract, as the model lacks explicit physical structure. These results highlight the advantage of SKANODE in simultaneously achieving high predictive accuracy and interpretable latent dynamics through its structured state-space formulation and symbolic discovery capability.
@@ 4.3 F-16 Aircraft
The last example expands to a real-world complex system. The study presented by [62] details the experimental data collected from a full-scale F-16 aircraft during a ground vibration test master class. The dataset is publicly available at https://data.4tu.nl/articles/_/12954911. To simulate real-world conditions, two dummy payloads were mounted on the wing tips of the aircraft. The setup includes 145 acceleration sensors distributed across the aircraft, under excitation of a shaker installed beneath the right wing to apply multisine excitations over a frequency range of 2 to 15 Hz. Data from three specific sensors were made public: one at the excitation point, one on the right wing near the nonlinear interface of interest, and one on the payload adjacent to the same interface. These measurements were taken at a sampling frequency of 400 Hz. The interfaces, consisting of T-shaped connectors on the payload side, slid through a rail attached to the wing side, were identified by preliminary investigation as primary sources of nonlinearity in the aircraft’s structural dynamics, particularly at the right-wing-to-payload interface.
@@
Building on these observations, SKANODE is extended by introducing an additional latent hysteretic state $a_{h}$, leading to the following structured formulation:
@@
where $h_{\theta}$ captures the hysteretic dynamics and $\alpha\in[0,1]$ weights the relative contributions of elastic stiffness and hysteretic components.
Through this extended structured state-space design, SKANODE is able to identify a symbolic governing equation, as presented in Figure 5. The resulting equation again aligns with the findings of [63], confirming that under low-level excitation, the system behavior is dominated by linear stiffness with mild hysteretic effects.
In comparison, the numerical baseline using SINDy fails to identify an equation incorporating the internal hysteretic mechanism. This is because it directly integrates the acceleration signals and seeks an explicit symbolic relation between acceleration, displacement, and velocity, without modeling any latent hysteresis effects. The identified equation from the numerical baseline, as given by
$a=0.0846-9.2480x-11.0607v+1.2781u+720.9876xv-12.2085x^{2}v,$
can reconstruct the observed accelerations to some extent but does not capture the true underlying system dynamics or reveal the implicit hysteretic variable.
Figure 6: Prediction error distribution for the F-16 acceleration response. SKANODE errors are more tightly clustered around zero and exhibit a distribution closer to normality compared with the deep learning baseline (ANODE) and the numerical SINDy baseline, indicating higher accuracy and more stable predictions.
Table 3: Model performance on F-16 acceleration prediction measured in MSE and SSIM. SKANODE outperforms both the deep learning baseline (ANODE) and the numerical baseline.
Model MSE SSIM
Numerical 0.0442 0.5868
ANODE 0.0456 0.5684
SKANODE 0.0343 0.6185
We further assess the model performance using both MSE and SSIM metrics. As reported in Table 3, the proposed SKANODE consistently outperforms the deep learning baseline (ANODE) as well as the numerical SINDy baseline in predicting accelerations. The distribution of prediction errors is revealed in Figure 6, where SKANODE exhibits errors that are more tightly clustered around zero and exhibit a distribution closer to normality compared with the baselines. This indicates that SKANODE not only achieves higher accuracy but also provides more stable and reliable predictions. Overall, this example underscores the capability of SKANODE to effectively capture complex, real-world nonlinear dynamics through the structured latent state design and interpretable symbolic equation discovery.
@@ 5 Conclusion
In this work, a novel framework, termed Structured Kolmogorov–Arnold Neural ODEs (SKANODE), is proposed, which integrates structured state-space modeling with symbolic equation discovery for learning interpretable dynamics from partially observed nonlinear systems. By embedding a physically meaningful latent state structure and leveraging the Kolmogorov–Arnold Network (KAN) as both a universal function approximator and symbolic learner, SKANODE bridges the gap between the expressive capacity of deep learning and the interpretability required for scientific modeling. This approach enables the direct recovery of physically meaningful latent states, such as displacements and velocities, while simultaneously extracting explicit symbolic governing equations directly from indirect measurements like accelerations.
Comprehensive experiments on multiple nonlinear dynamical systems demonstrate that SKANODE outperforms existing NODE variants in both predictive accuracy and interpretability. These results highlight the importance of combining inductive structural biases with symbolic discovery capabilities to guide the learning process toward solutions that are physically consistent, robust, and interpretable. Specifically, SKANODE accurately recovers governing equations that align with ground truth on classical nonlinear benchmarks, including the Duffing and Van der Pol oscillators. In the real-world F-16 aircraft system, SKANODE identifies complex hysteretic behavior and derives an interpretable symbolic equation that captures this nonlinear phenomenon.
The ability to uncover governing equations from indirect observations makes SKANODE particularly well-suited for scientific and engineering domains where full-state measurements are often impractical, and model transparency is critical for trustworthy decision-making. Future work will explore the integration of domain-specific physics priors and constraints within the symbolic extraction process.
Another important direction concerns robustness to real-world sensing imperfections such as sensor drift, measurement bias, and colored noise. While SKANODE has been validated under varying white noise levels on simulated systems such as the Duffing oscillator, future work may further investigate SKANODE’s ability to extract latent dynamics from measurements affected by structured or colored noise. One promising extension is to incorporate a decoupled observation model, where the output mapping is trained with the dynamics to account for sensing distortions. Such modeling flexibility would allow SKANODE to adapt to more challenging real-world sensing environments.
@@ Appendix A SKANODE Recovers True System Dynamics
In this section, it is theoretically demonstrated that, under certain conditions and assuming idealized loss minimization in the theoretical setting, the proposed Structured Kolmogorov–Arnold Neural ODE (SKANODE) framework is capable of exactly recovering the true governing dynamics.